OpenStudy (anonymous):

Evaluate the expression below when x = 3. -2x - 8 -(-3x + 8)

OpenStudy (anonymous):

plug 3 in for x. -2(3)-8-(-3(3)+8) now solve.

OpenStudy (anonymous):

-2(3)-8-(-3(3)+8) =-6-8-(-9+8) =-6-8-(-1) =-14+1 =-13